### v1

@@ 3.4 Radial gradients: LTE vs NLTE
We measured the chemical abundances under the LTE assumption and performed the logarithmic fit of the LTE abundance gradients. Fig. 15 reveals that for almost all the elements the NLTE slopes of the logarithmic fit are shallower. Only S and Cu show a similar slope, which suggests negligible NLTE effects for the lines employed for these two elements. This is expected for S, since we used an S I triplet that is weakly prone to NLTE effects (duffau2017gaia). NLTE have a minor impact on Cu I lines in stars with dex. shi18 found a mean abundance increase of about +0.07 dex when applying NLTE corrections, a result previously suggested by yan16.
@@
The histograms in Fig. 16 illustrate the comparison of the NLTE and LTE abundances where, on average, lower NLTE abundances for O (-0.74 0.21 dex), Na (-0.110.08 dex), Al (-0.070.10 dex), Si (-0.080.15 dex) and Cu (-0.130.17 dex), and more centered distributions for Mg (-0.030.08 dex), S (-0.020.10 dex), Ca (0.060.09 dex) and Mn (0.020.14 dex). Our results indicate that the oxygen lines used in this study (OI 777 nm triplet) are strongly affected by NLTE, consistent with amarsi16, who reported corrections of about 0.6 dex or higher at solar metallicities. For the other elements, the NLTE corrections are more moderate and agree with the findings of amarsi2020galah. In appendix D we further discuss the NLTE abundances for Mn and O.
@@
In Fig. 17, we compare the slopes of our chemical gradients with those reported in previous studies. A general agreement is observed in the inner disk for most elements, although a slight underabundance is found for sulfur and copper when comparing the GPR profiles to the linear trends from other works. Moreover, the GPR profiles provide additional information on the behavior of Na, Mg, Al, S, Ca, Ti, Mn and Cu, revealing a shoulder at the solar circle similar to that obtained in the iron GPR profile. The GPR profiles also suggest a change of slope around 14-16 kpc, resulting in a flatter gradient in the outer disk compared to the linear trends of previous studies, particularly for Na, Mg, Al, Si, S, Ca, Ti, Mn and Cu.
@@
In Fig. 18, we emphasize the agreement with the da2023oxygen logarithmic profiles for oxygen and sulfur. According to da2023oxygen, the radial gradient of sulfur is steeper than that of iron, and sulfur is on average under-abundant compared to oxygen. This discrepancy points to the need for a revision of chemical evolution models, particularly regarding sulfur yields from massive stars, in order to reproduce observed trends in the outer Galaxy. However, our results confirm the general behavior of sulfur but show a profile that is more consistent with iron and the other elements, rather than reproducing the steeper trend reported by da2023oxygen.
@@ 3.5 Kinematic properties
The kinematic properties of CCs provide fundamental insights to constrain their true nature of thin disk stellar population. Their orbits were integrated by using the galpy code, a Python library dedicated to Galactic dynamics (bovy15). Positional parameters (RA, DEC, distance) and proper motions are mainly based on Gaia DR3. Individual distances have been discussed in more detail in § 2.1, while radial velocities are mainly based on current spectroscopic measurements. The galactic potential employed is MWPotential2014.
@@
To improve the sampling across the four quadrants, the current sample was complemented with CCs analysed by luck18 and by Trentin24B for which homogeneous elemental abundances are available. We did not include the 16 CCs located at 3–5.6 kpc and analyzed by using high resolution NIR spectra by matsunaga23, because of possible systematics between NIR and optical spectroscopy. Moreover, the quoted samples have a sizable number of objects in common with the current sample and they have been adopted to move their chemical abundances into our metallicity scale.
@@
Fig. 19 shows nine different kinematic planes with the orbital properties of CCs compared with a sample of Type II Cepheids (TIICs). These planes are based on kinematic diagnostics that are quite useful to characterize Galactic stellar populations, and in particular to identify stars that are associated to the the different Galactic components (lane22; bonifacio24; Bono et al. submitted). According to a Galactocentric cylindrical frame, , and are the radial, the tangential and the vertical velocity, while and are the angular momentum component along the Galactic Z-axis and the total orbital energy. , , and are the radial, the azimuthal, and the vertical actions, while is the total action, defined as . is the normalized angular momentum (circularity), defined as where is the angular momentum of a circular orbit with the same maximum binding energy. Finally, denotes the maximum height along the orbit above the Galactic plane. The bulk of the kinematic sample has cold orbits typical of thin disk stars, as stated by the coherent distribution in the 9 planes. In particular, in the Toomgre diagram (panel a) our sample is grouped in the bottom-right corner, characterized by and low perpendicular velocities, typical of the thin disk component. On the other hand, TIICs cover a broad range in transversal velocity and a fraction of them also attain negative transversal velocities suggesting that they are on retrograde orbits. The stark difference between the two samples is fully supported by the Lindblad plane (panel c) in which CCs are distributed along the tiny region typical of thin disk stars, while TIICs move from the region typical of hot stellar components with radial orbits and random motions. The same outcome applies to the action-diamond diagram (panel f), since CCs are grouped in the right corner, since the major contribution to the action is given by their . Furthermore, CCs show high circularities (, which means almost circular orbits and strong rotation), low eccentricities and low values (panels g, h, i). The TIICs display a more complex behavior, since they move from a warm stellar component more typical of thick disk stars to hot stellar components and to retrograde orbits. A more quantitative analysis will be provided in a forthcoming paper (Nunnari et al. in preparation).
@@
The left panel of Fig. 20 shows the radial distribution of CCs projected onto the Galactic plane. The symbols are color-coded according to the azimuth angle . We note that stars with typically reach smaller Galactocentric radii than those at . This trend may arise from the intrinsic spatial distribution of stars in the sample (for example, due to the presence of spiral arms), from the selection function of our dataset, or from a combination of both effects. Data plotted in the central and right panel of the same figure display that for Galactocentric distances larger than 11-12 kpc the height above the Galactic plane steadily decreases. This finding is expected, since the variation is mainly caused by the Galactic warp. Specifically, it is known from previous observations that the Galactic midplane is bent downwards (i.e. Z¡0) in the portion of the disk covered by our dataset. The reader interested in a more detailed discussion concerning the Galactic warp is referred to chen19, skowron19, lemasle22 and poggio25.
@@
It is worth mentioning that the change in the slope of the radial metallicity gradient occurs at Galactocentric radii where the Galactic warp begins to become significant (i.e. approximately 11-12 kpc), tentatively suggesting that the two aspects might be related. It is important to note, however, that the spiral arms and bar-driven radial mixing can also potentially generate variations in the radial metallicity gradient (see e.g., buder25).
@@
The orbital properties (radial velocity and circularity) of these objects fully support their association to the thin disk. However, more photometric and spectroscopic data are required to address this issue on a more quantitative basis.
@@ 4 Summary and conclusions
Previous studies available in the literature have modeled the chemical gradients of the thin disk, by assuming either a linear (luck18; genovali2014fine; Trentin24B) or a logarithmic (da2023oxygen) function. The fits based on a linear function overestimate the observed gradient in the inner disk ( kpc) and underestimate the observed gradient in the outer disk ( kpc). The difference with previous investigations, based on a linear fit, is mainly due to the larger range in Galactocentric distances covered by the current sample. This means that we are dealing with 22 CCs with smaller than 6 kpc and 33 CCs with a larger than 15 kpc666The distance of the Cepheid ASAS J062939-1840.5 was estimated by using a mid-infrared (W1-band) PL relation and we found kpc. The distance of this object was also estimated by trentin23 using period-Wesenheit relations and they found kpc..
@@
A flattening of the abundance gradient in the outer disk has been reported by several studies (e.g., luck18; kovtyukh22; da2023oxygen). A similar behavior has also been observed in open clusters (magrini23; otto25), where a bilinear fitting approach has been adopted. Nevertheless, the AIC test performed by otto25 indicated that introducing two additional parameters in a bilinear model does not significantly improve the fit compared to a simple linear relation. In any case, such modeling inherently implies a break at a certain Galactocentric radius, raising a theoretical question about the physical origin of this discontinuity. For our sample, the best-fitting bilinear model applied to the iron gradient (Fig. 5) suggests a break, or ”knee”, at kpc. This model provides a better fit than the linear one, as indicated by smaller residuals and a lower AIC score. However, when the same bilinear approach is applied to the other elements, the inferred knee positions vary considerably, ranging from 8 to 15 kpc. The bilinear model yields the lowest AIC score only for Mg compared to alternative models (Fig. 11), although in this case the knee occurs at kpc, a value inconsistent with the other elements and with previous studies.
@@
In da2023oxygen, a large sample of 356 CCs, including four at kpc, was analyzed using a log–log relation for Fe, O, and S. In Fig. 5, we applied the same functional form to fit the iron gradient. As expected, this approach yields lower mean squared error, chi-squared, and AIC values compared to the linear and bilinear models, except in the case of Mg, as noted above. Furthermore, as shown in Fig. 7, the slope we obtain differs from that reported in da2023oxygen, likely due to differences in spectral analysis, such as the line list and the use of NLTE grids for estimating atmospheric parameters, and to the increased number of Cepheids located in the outskirt of the disk. Overall, the log-log fit appears to be the most reasonable among the three analytical models, although sodium and magnesium are better modeled by the bilinear fit (Fig. 11), suggesting that a single analytical function may not fully describe all gradients. For this reason, we also employed Gaussian Process Regression, which does not require any predetermined functional form. This modeling confirms that the radial abundance gradients in the thin disk generally flatten with increasing Galactocentric distance across all studied elements.
@@
The flattening of the gradients is a feature that extends to extragalactic studies. Works on other disk galaxies have shown that, although the inner regions show a negative gradient, the radial distribution of metallicity flattens to a virtually constant value beyond the isophote radius. This behavior has been found in the disk galaxy M83 (bresolin09, bresolin16), and subsequently confirmed in NGC 1512 and NGC 3621 (bresolin12; kudritki14), where the oxygen abundance remains homogeneous and approaches an almost constant value in the outer regions. These observations suggest that flattening is a common feature of the metallicity gradient in spiral galaxy disks.
@@
The current metallicity of CCs can be compared with that of B-stars located in the solar neighborhood. According to nieva12, early B-type stars within 500 pc show a mean iron abundance of 7.520.03 dex, consistent with the solar value. This result is based on a NLTE spectral analysis of 29 early B-type stars. In the present work, the mean iron abundance of the 11 CCs located within 500 pc of the Sun is 7.490.13 dex, in perfect agreement with nieva12. Differently, the 93 CCs located at the solar circle (i.e. all Cepheids with kpc), exhibit a slightly iron underabundance of 7.420.10 dex, but still consistent within 1.
@@
Our results for chemical abundance ratios as a function of reveal that [X/Fe] ratios remain approximately constant across the entire thin disk. The near-flat distributions, with average values within 0.10 dex and minimal slopes, indicate that most elemental abundances scale closely with iron throughout the Galactic disk. This consistency suggests that these elements trace the gas distribution in the thin disk in a coherent manner. Slight overabundances in Al and Na, underabundance in Cu, and the small but measurable slopes for Mn, and Cu point to subtle variations, yet they do not compromise the overall trend of roughly constant abundance ratios. The negative slopes observed for Mn and Cu are in agreement with previous findings for iron-peak elements (otto25).
@@
Chemical gradients from previous studies, such as genovali2014fine and da2023oxygen, indicate that the outer disk exhibits a higher dispersion. This effect can be partially attributed to the Galactic warp (kerr57; oort58; lemasle22; poggio25), a large-scale distortion of the disk that affects the estimated Galactocentric distances. In regions where the warp is pronounced, Cepheids appear closer in Galactocentric distance than they would in a non-warped disk. However, lemasle22 noted that the dispersion introduced by the warp is too small to fully account for the larger spread around the mean metallicity gradient reported by genovali2014fine and da2023oxygen. Instead, this increased dispersion likely reflects larger uncertainties associated with distant stars, whose spectra typically have lower signal-to-noise ratio. In our study, the gradients do not show a significant increase in dispersion (Fig. 10), in agreement with lemasle22. It is worth noting that our sample contains fewer Cepheids in the outer disk ( kpc), where the warp is stronger, than in the inner regions. Nevertheless, the trend visible in Fig. 20 appears to be a direct consequence of the Galactic warp. Similar trends have also been observed using larger kinematic samples of Cepheids (lemasle22, Fig. A.3; poggio25).
@@
Results for the sample of the youngest open clusters (OCs) from otto25 overlaps with that of Cepheids. Sodium is an exception, since Cepheids are on average overabundant in this element. It was suggested by takeda13 and confirmed by dasilva15 in a comparison between giant and dwarf field stars, that this overabundance is mainly due to stellar evolution effects: the first dredge-up brings some sodium from the interior to the atmosphere, altering its initial chemical composition (ventura13; lagarde21). In contrast, the chemical abundances of OCs have been measured in both giant and dwarf stars, with the latter still unaffected by significant mixing processes. Comparisons with other stellar populations, such as open clusters, generally reveal similar radial trends for iron-peak and alpha elements, further supporting the homogeneity of chemical enrichment in the thin disk, at least in the region between 5 and 14 kpc. These findings are consistent with previous studies (magrini23; yong12; otto25).
@@
The chemical abundances of other elements, such as heavy elements produced by the s- and r-processes, have yet to be determined. As discussed in Trentin24B, some elements-such as Scandium, Zinc, and Zirconium-show a bifurcation at the low-metallicity tail. This anomalous behavior may arise from variations in spectral line analyses or from the presence of distinct stellar populations associated with different spiral arms. Since our sample includes Cepheids located in various spiral arms, a future study focusing on these elements could help clarify this issue. However, NLTE grids for these elements are not yet available in PySME.
In the inner disk, for kpc, a double sequence can appear, one that follows the logarithmic profile and the other showing a change in the slope. Note that a significant decrease in the metallicity of the young stellar population, confined to the central/inner regions, has been found in barred spiral galaxies such as NGC 1365 and M83 (sextl24, sextl25). Furthermore, andrievsky16 discussed a change in the slope at kpc, suggesting that it might be caused by a reduced star formation rate compared to the outer disk regions. To account for such a change of slope in the radial gradients, it has also been suggested that the inner disk regions have also been affected by dilution effects caused by the infall of more metal-poor gas (genovali15, and references therein). Further extensive sampling at small is crucial to investigate this feature of the gradient and its implications for the chemical enrichment of the Galactic bar.
The [X/H] gradients exhibit similar slopes among the studied elements, and the [X/Fe] ratios remain approximately constant across the entire range of explored. This behavior may indicate that these elements could serve as potential tracers of the underlying gas density profile. However, further validation is required to confirm these trends, and it remains to be established whether such a result can be extended to the disks of external galaxies.
@@
In conclusion, our results suggest that a single analytic function is not sufficient to model all the gradients. For this reason, we provided the behavior of each element gradient with a GPR modeling. However, when using an analytical function, a log-log profile should be preferred. Our results confirm that the radial abundance gradients in the thin disk flatten with radius across all studied elements. This is consistent with theoretical expectations of inside-out disk formation and continuous chemical enrichment, where inner regions exhibit higher metallicity and alpha-enhancements due to more rapid star formation. Nevertheless, the outer disk is still poorly sampled to give a definitive conclusion.
The adoption of NLTE modeling throughout, from atmospheric parameters to elemental abundances, represents an important step forward relative to prior studies that often combined LTE atmospheric parameters with NLTE corrections. The method employed in this study could be extended to data that come from large spectroscopic surveys (4MOST777https://www.4most.eu/cms/home/, dejong16) and from new instruments (MOONS@VLT888https://vltmoons.org/, cirasuolo12; MAVIS@VLT999https://mavis-ao.org/mavis/, mcdermid20), which will allow to increase sample sizes and extend spatial coverage, allowing for finer chemical gradient mapping, especially in regions that we are now poorly sampling such as the inner and outer disk, and improved constraints on Galactic formation models.

### v2

@@ 3.4 Radial gradients: LTE versus NLTE
We measured the chemical abundances under the LTE assumption and performed the logarithmic fit of the LTE abundance gradients. Figure 15 reveals that for almost all the elements the NLTE slopes of the logarithmic fit are shallower. Only S and Cu show a similar slope, which suggests negligible NLTE effects for the lines employed for these two elements. This is expected for S, since we used an S I triplet that is weakly prone to NLTE effects (Duffau et al. 2017). NLTE have a minor impact on Cu I lines in stars with dex. Shi et al. (2018) found a mean abundance increase of about +0.07 dex when applying NLTE corrections, a result previously suggested by Yan et al. (2016).
@@
The histograms in Fig. 16 illustrate the comparison of the NLTE and LTE abundances where, on average, lower NLTE abundances for O ( dex), Na ( dex), Al ( dex), Si ( dex) and Cu ( dex), and more centred distributions for Mg ( dex), S ( dex), Ca ( dex) and Mn ( dex). Our results indicate that the oxygen lines used in this study (OI 777 nm triplet) are strongly affected by NLTE, consistent with Amarsi et al. (2016), who reported corrections of about 0.6 dex or higher at solar metallicities. For the other elements, the NLTE corrections are more moderate and agree with the findings of Amarsi et al. (2020).
We emphasise that in our analysis, the derivation of abundances is not performed at fixed atmospheric parameters: each run (LTE or NLTE) adopts a consistent set of atmospheric parameters optimised for that assumption. As a consequence, the comparison of abundances between NLTE and LTE cannot be attributed solely to the direct NLTE line corrections for a given element, since the underlying atmospheric parameters themselves also differ. In appendix D we further discuss the NLTE abundances for Mn and O.
@@
In Fig. 17, we compare the slopes of our chemical gradients with those reported in previous studies. A general agreement is observed in the inner disk for most elements, although a slight underabundance is found for sulfur and copper when comparing the GPR profiles to the linear trends from other works. Moreover, the GPR profiles provide additional information on the behaviour of Na, Mg, Al, S, Ca, Ti, Mn and Cu, revealing a shoulder at the solar circle similar to that obtained in the iron GPR profile. The GPR profiles also suggest a change of slope around 14-16 kpc, resulting in a flatter gradient in the outer disk compared to the linear trends of previous studies, particularly for Na, Mg, Al, Si, S, Ca, Ti, Mn and Cu.
@@
In Fig. 18, we emphasize the agreement with the dS23 logarithmic profiles for oxygen and sulfur. According to dS23, the radial gradient of sulfur is steeper than that of iron, and sulfur is on average under-abundant compared to oxygen. This discrepancy points to the need for a revision of chemical evolution models, particularly regarding sulfur yields from massive stars, in order to reproduce observed trends in the outer Galaxy. However, our results confirm the general behaviour of sulfur but show a profile that is more consistent with iron and the other elements, rather than reproducing the steeper trend reported by dS23.
@@ 3.5 Kinematic properties
The kinematic properties of CCs provide fundamental insights to constrain their true nature of thin disk stellar population. Their orbits were integrated by using the galpy code, a Python library dedicated to Galactic dynamics (Bovy 2015). Positional parameters (RA, DEC, distance) and proper motions are mainly based on Gaia DR3. Individual distances have been discussed in more detail in Sect. 2.1, while radial velocities are mainly based on current spectroscopic measurements. The galactic potential employed is MWPotential2014.
@@
To improve the sampling across the four quadrants, the current sample was complemented with CCs analysed by Luck (2018) and by Trentin et al. (2024a) for which homogeneous elemental abundances are available. We did not include the 16 CCs located at 3-5.6 kpc and analysed by using high resolution NIR spectra by Matsunaga et al. (2023), because of possible systematics between NIR and optical spectroscopy. Moreover, the quoted samples have a sizable number of objects in common with the current sample and they have been adopted to move their chemical abundances into our metallicity scale.
@@
Figure 19 shows nine different kinematic planes with the orbital properties of CCs compared with a sample of Type II Cepheids (TIICs). These planes are based on kinematic diagnostics that are quite useful to characterize Galactic stellar populations, and in particular to identify stars that are associated with the different Galactic components (Lane et al. 2022; Bonifacio et al. 2024; Bono et al. 2025). According to a Galactocentric cylindrical frame, , and are the radial, the tangential and the vertical velocity, while and are the angular momentum component along the Galactic Z-axis and the total orbital energy. , , and are the radial, the azimuthal, and the vertical actions, while is the total action, defined as . is the normalized angular momentum (circularity), defined as where is the angular momentum of a circular orbit with the same maximum binding energy. Finally, denotes the maximum height along the orbit above the Galactic plane. The bulk of the kinematic sample has cold orbits typical of thin disk stars, as stated by the coherent distribution in the 9 planes. In particular, in the Toomgre diagram (panel a) our sample is grouped in the bottom-right corner, characterized by and low perpendicular velocities, typical of the thin disk component. On the other hand, TIICs cover a broad range in transversal velocity and a fraction of them also attain negative transversal velocities suggesting that they are on retrograde orbits. The stark difference between the two samples is fully supported by the Lindblad plane (panel c) in which CCs are distributed along the tiny region typical of thin disk stars, while TIICs move from the region typical of hot stellar components with radial orbits and random motions. The same outcome applies to the action-diamond diagram (panel f), since CCs are grouped in the right corner, since the major contribution to the action is given by their . Furthermore, CCs show high circularities (, which means almost circular orbits and strong rotation), low eccentricities and low values (panels g, h, i). The TIICs display a more complex behaviour, since they move from a warm stellar component more typical of thick disk stars to hot stellar components and to retrograde orbits. A more quantitative analysis will be provided in a forthcoming paper (Nunnari et al. in preparation).
@@
The left panel of Fig. 20 shows the radial distribution of CCs projected onto the Galactic plane. The symbols are colour-coded according to the azimuth angle . We note that stars with typically reach smaller Galactocentric radii than those at . This trend may arise from the intrinsic spatial distribution of stars in the sample (for example, due to the presence of spiral arms), from the selection function of our dataset, or from a combination of both effects. Data plotted in the central and right panel of the same figure display that for Galactocentric distances larger than 11-12 kpc the height above the Galactic plane steadily decreases. This finding is expected, since the variation is mainly caused by the Galactic warp. Specifically, it is known from previous observations that the Galactic midplane is bent downwards (i.e. Z¡0) in the portion of the disk covered by our dataset. The reader interested in a more detailed discussion concerning the Galactic warp is referred to Chen et al. (2019), Skowron et al. (2019), Lemasle et al. (2022) and Poggio et al. (2025).
@@
It is worth mentioning that the change in the slope of the radial metallicity gradient occurs at Galactocentric radii where the Galactic warp begins to become significant (i.e. approximately 11-12 kpc), tentatively suggesting that the two aspects might be related. It is important to note, however, that the spiral arms and bar-driven radial mixing can also potentially generate variations in the radial metallicity gradient (see e.g., Buder et al. 2025).
@@
The orbital properties (radial velocity and circularity) of these objects fully support their association with the thin disk. However, more photometric and spectroscopic data are required to address this issue on a more quantitative basis.
@@ 4 Summary and conclusions
Previous studies available in the literature have modeled the chemical gradients of the thin disk, by assuming either a linear (Luck 2018; Genovali et al. 2014; Trentin et al. 2024a) or a logarithmic (da Silva et al. 2023) function. The fits based on a linear function overestimate the observed gradient in the inner disk ( kpc) and underestimate the observed gradient in the outer disk ( kpc). The difference with previous investigations, based on a linear fit, is mainly due to the larger range in Galactocentric distances covered by the current sample. This means that we are dealing with 22 CCs with smaller than 6 kpc and 33 CCs with a larger than 15 kpc666The distance of the Cepheid ASAS J062939-1840.5 was estimated by using a mid-infrared (W1-band) PL relation and we found kpc. The distance of this object was also estimated by Trentin et al. (2023) using period-Wesenheit relations and they found kpc..
@@
A flattening of the abundance gradient in the outer disk has been reported by several studies (e.g., Luck 2018; Kovtyukh et al. 2022; da Silva et al. 2023). A similar behaviour has also been observed in open clusters (Magrini et al. 2023; Otto et al. 2025), where a bilinear fitting approach has been adopted. Nevertheless, the AIC test performed by Otto et al. (2025) indicated that introducing two additional parameters in a bilinear model does not significantly improve the fit compared to a simple linear relation. In any case, such modelling inherently implies a break at a certain Galactocentric radius, raising a theoretical question about the physical origin of this discontinuity. For our sample, the best-fitting bilinear model applied to the iron gradient (Fig. 5) suggests a break, or ”knee”, at kpc. This model provides a better fit than the linear one, as indicated by smaller residuals and a lower AIC score. However, when the same bilinear approach is applied to the other elements, the inferred knee positions vary considerably, ranging from 8 to 15 kpc. The bilinear model yields the lowest AIC score only for Mg compared to alternative models (Fig. 11), although in this case the knee occurs at kpc, a value inconsistent with the other elements and with previous studies.
@@
In dS23, a large sample of 356 CCs, including four at kpc, was analysed using a log–log relation for Fe, O, and S. In Fig. 5, we applied the same functional form to fit the iron gradient. As expected, this approach yields lower mean squared error, chi-squared, and AIC values compared to the linear and bilinear models, except in the case of Mg, as noted above. Furthermore, as shown in Fig. 7, the slope we obtain differs from that reported in dS23, likely due to differences in spectral analysis, such as the line list and the use of NLTE grids for estimating atmospheric parameters, and to the increased number of Cepheids located in the outskirt of the disk. Overall, the log-log fit appears to be the most reasonable among the three analytical models, although sodium and magnesium are better modeled by the bilinear fit (Fig. 11), suggesting that a single analytical function may not fully describe all gradients. For this reason, we also employed Gaussian Process Regression, which does not require any predetermined functional form. This modelling confirms that the radial abundance gradients in the thin disk generally flatten with increasing Galactocentric distance across all studied elements.
@@
The flattening of the gradients is a feature that extends to extragalactic studies. Works on other disk galaxies have shown that, although the inner regions show a negative gradient, the radial distribution of metallicity flattens to a virtually constant value beyond the isophote radius. This behaviour has been found in the disk galaxy M83 (Bresolin et al. 2009, 2016), and subsequently confirmed in NGC 1512 and NGC 3621 (Bresolin et al. 2012; Kudritzki et al. 2014), where the oxygen abundance remains homogeneous and approaches an almost constant value in the outer regions. These observations suggest that flattening is a common feature of the metallicity gradient in spiral galaxy disks.
@@
The current metallicity of CCs can be compared with that of B-stars located in the solar neighborhood. According to Nieva and Przybilla (2012), early B-type stars within 500 pc show a mean iron abundance of 7.520.03 dex, consistent with the solar value. This result is based on a NLTE spectral analysis of 29 early B-type stars. In the present work, the mean iron abundance of the 11 CCs located within 500 pc of the Sun is 7.490.13 dex, in perfect agreement with Nieva and Przybilla (2012). Differently, the 93 CCs located at the solar circle (i.e. all Cepheids with kpc), exhibit a slightly iron underabundance of 7.420.10 dex, but still consistent within 1.
@@
Our results for chemical abundance ratios as a function of reveal that [X/Fe] ratios remain approximately constant across the entire thin disk. The near-flat distributions, with average values within 0.10 dex and minimal slopes, indicate that most elemental abundances scale closely with iron throughout the Galactic disk. This consistency suggests that these elements trace the gas distribution in the thin disk in a coherent manner. Slight overabundances in Al and Na, underabundance in Cu, and the small but measurable slopes for Mn, and Cu point to subtle variations, yet they do not compromise the overall trend of roughly constant abundance ratios. The negative slopes observed for Mn and Cu are in agreement with previous findings for iron-peak elements (Otto et al. 2025).
@@
Chemical gradients from previous studies, such as Genovali et al. (2014) and dS23, indicate that the outer disk exhibits a higher dispersion. This effect can be partially attributed to the Galactic warp (Kerr et al. 1957; Oort et al. 1958; Lemasle et al. 2022; Poggio et al. 2025), a large-scale distortion of the disk that affects the estimated Galactocentric distances. In regions where the warp is pronounced, Cepheids appear closer in Galactocentric distance than they would in a non-warped disk. However, Lemasle et al. (2022) noted that the dispersion introduced by the warp is too small to fully account for the larger spread around the mean metallicity gradient reported by Genovali et al. (2014) and da Silva et al. (2023). Instead, this increased dispersion likely reflects larger uncertainties associated with distant stars, whose spectra typically have lower signal-to-noise ratio. In our study, the gradients do not show a significant increase in dispersion (Fig. 10), in agreement with Lemasle et al. (2022). It is worth noting that our sample contains fewer Cepheids in the outer disk ( kpc), where the warp is stronger, than in the inner regions. Nevertheless, the trend visible in Fig. 20 appears to be a direct consequence of the Galactic warp. Similar trends have also been observed using larger kinematic samples of Cepheids (Lemasle et al. 2022, Fig. A.3; Poggio et al. 2025).
@@
Results for the sample of the youngest open clusters (OCs) from Otto et al. (2025, age ¡ 400 Myrs) overlaps with that of Cepheids. Sodium is an exception, since Cepheids are on average overabundant in this element. It was suggested by Takeda et al. (2013) and confirmed by da Silva et al. (2015) in a comparison between giant and dwarf field stars, that this overabundance is mainly due to stellar evolution effects: the first dredge-up brings some sodium from the interior to the atmosphere, altering its initial chemical composition (Ventura et al. 2013; Lagarde et al. 2012). In contrast, the chemical abundances of OCs have been measured in both giant and dwarf stars, with the latter still unaffected by significant mixing processes. Comparisons with other stellar populations, such as open clusters, generally reveal similar radial trends for iron-peak and alpha elements, further supporting the homogeneity of chemical enrichment in the thin disk, at least in the region between 5 and 14 kpc. These findings are consistent with previous studies (Magrini et al. 2023; Yong et al. 2012; Otto et al. 2025).
@@
The chemical abundances of other elements, such as heavy elements produced by the s- and r-processes, have yet to be determined. As discussed in Trentin et al. (2024a), some elements such as scandium, zinc, and zirconium show a bifurcation at the low-metallicity tail. This anomalous behaviour may arise from variations in spectral line analyses or from the presence of distinct stellar populations associated with different spiral arms. Since our sample includes Cepheids located in various spiral arms, a future study focusing on these elements could help clarify this issue. However, NLTE grids for these elements are not yet available in PySME.
In the inner disk, for kpc, a double sequence can appear, one that follows the logarithmic profile and the other showing a change in the slope. Note that a significant decrease in the metallicity of the young stellar population, confined to the central and inner regions, has been found in barred spiral galaxies such as NGC 1365 and M83 (Sextl et al. 2024, 2025). Furthermore, Andrievsky et al. (2016) discussed a change in the slope at kpc, suggesting that it might be caused by a reduced star formation rate compared to the outer disk regions. To account for such a change of slope in the radial gradients, it has also been suggested that the inner disk regions have also been affected by dilution effects caused by the infall of more metal-poor gas (Genovali et al. 2015, and references therein). Further extensive sampling at small is crucial to investigate this feature of the gradient and its implications for the chemical enrichment of the Galactic bar.
The [X/H] gradients exhibit similar slopes among the studied elements, and the [X/Fe] ratios remain approximately constant across the entire range of explored. This behaviour may indicate that these elements could serve as potential tracers of the underlying gas density profile. However, further validation is required to confirm these trends, and it remains to be established whether such a result can be extended to the disks of external galaxies.
@@
In conclusion, our results suggest that a single analytic function is not sufficient to model all the gradients. For this reason, here we provide the behaviour of each element gradient with a GPR modelling. However, when using an analytical function, a log-log profile should be preferred. Our results confirm that the radial abundance gradients in the thin disk flatten with radius across all studied elements. This is consistent with theoretical expectations of inside-out disk formation and continuous chemical enrichment, where inner regions exhibit higher metallicity and alpha-enhancements due to more rapid star formation. Nevertheless, the outer disk is still too poorly sampled to give a definitive conclusion.
The adoption of NLTE modelling throughout, from atmospheric parameters to elemental abundances, represents an important step forward relative to prior studies that often combined LTE atmospheric parameters with NLTE corrections. The method employed in this study could be extended to data that come from large spectroscopic surveys (4MOST777https://www.4most.eu/cms/home/, de Jong et al. 2016) and from new instruments (MOONS888https://vltmoons.org/ at VLT, Cirasuolo et al. 2012; MAVIS999https://mavis-ao.org/mavis/ at VLT, Content et al. 2022). This will allow for an increase in sample sizes and extension of the spatial coverage, enabling finer chemical gradient mapping, especially in regions now poorly sampled such as the inner and outer disk, along with setting improved constraints on Galactic formation models.
5 Data availability
Tables 11, 12, 13, 14 and 15 are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.